# Leveraging Data Augmentation and Siamese Learning for Predictive Process Monitoring

### v1

@@ 1 Introduction
Due to the increasing digitization of business operations, Predictive Process Monitoring (PPM) has become a critical area of research within process mining. Two of its core tasks, next activity prediction and final outcome prediction, have received substantial attention due to their practical relevance in optimizing workflows, reducing delays, and improving resource utilization [20]. While earlier works utilized symbolic models such as Petri nets and transition systems, recent advancements in deep learning, especially sequence models such as Long-Short Term Memory (LSTM), Gated Recurrent Unit and attention-based Transformers, have significantly improved the predictive performance of such models. However, a persistent challenge in deep learning and inherently in its application to PPM, is the limited size of available labeled data [3]. Despite their high complexity, business processes are typically logged with limited trace variation and often result in datasets containing redundant or highly similar samples. These limitations affect the generalizability of predictive models, especially in scenarios with infrequent behaviors or class imbalance. In machine learning, a common remedy for such constraints is data augmentation: the process of artificially generating additional training data to enhance diversity and robustness. Data augmentation has been extensively explored in domains such as computer vision and natural language processing (NLP) [21]. In these fields, augmentations are informed by domain knowledge to preserve the semantic integrity of the data. For example, visual enhancements such as flipping or cropping retain object identity, while textual methods such as synonym replacement, word deletion or insertion aim to maintain grammatical and contextual coherence [22]. The core principle in these domains is that augmentations should be semantically valid and task-relevant. In contrast, data augmentation for process mining remains considerably underexplored. A recent work introduces model-agnostic augmentation for event logs [12], drawing inspiration from Easy Data Augmentation (EDA) transformations to randomly insert, delete, or replace activities [22]. However, this method applies transformations uniformly at random, without incorporating control-flow semantics or behavioral statistics. This introduces several critical shortcomings. Random modifications may produce traces that contradict the logic of the process model, e.g. placing an approval activity before a request or breaking mandatory activity sequences, violating the process semantics (cf. Fig. 1).
Without incorporation of statistically frequent patterns in the process, augmented samples may not reflect realistic decision points or execution branches, thus limiting their relevance for generalization. Despite the importance of preserving control-flow behavior in PPM, no prior work has systematically addressed these limitations by designing statistically grounded augmentation techniques that align with the dynamics and semantics of business processes (cf. Section 2). This paper seeks to fill this gap by proposing a principled approach to data augmentation in PPM. We introduce three novel transformation techniques, named StatisticalInsertion, StatisticalDeletion and StatisticalReplacement. These techniques use frequency-based mining of event logs to generate meaningful variations of trace data. They explicitly capture frequent direct followers, intermediate activity sequences and XOR-splits to ensure that the augmented traces maintain control-flow while introducing variability.
@@
In addition to augmentation, we also explore the underutilized potential of self-supervised learning (SSL) in PPM. While most existing approaches rely on supervised learning, this requires labeled data (e.g. next activity or final outcome annotations) for each trace. Practically, such amount of labels may not be sufficient to train a model. In fact, some event logs may be imbalanced, resulting in, for example, an insufficient number of examples for a given outcome, e.g. a given machinary breaks down. This issue also applies to the next-activity prediction task, where the occurrence of certain activities may be rare, leading to a very limited number of events with the corresponding label. SSL addresses this by learning from the data itself without requiring manual labels. It constructs surrogate tasks, such as aligning different augmented views of the same input to learn robust task-agnostic representations. Although SSL has seen great success in the computer vision domain [4, 10, 5], it is underexplored in PPM. In this work, we adopt a Siamese learning strategy based on Bootstrap Your Own Latent (BYOL) [10], where augmented trace prefixes are encoded to similar latent representations. This setup avoids the need for negative samples or large batch sizes [10], making it well-suited for structured, small-scale process logs. Combined with our statistically grounded augmentations, this enables high-quality representation for downstream tasks, even with limited labeled data.
@@ 2.0.1 Data Augmentation Techniques
Data augmentation is a proven strategy to increase training data and improve model generalization. In NLP, a well-known method is EDA [22], which introduces random augmentations such as random insertion, deletion and replacement of words. Other techniques such as syntax-aware transformations [8] and generative approaches such as InsNet [15] inject linguistic diversity into text classification and translation tasks. In computer vision, simple pixel-level augmentations and mixing techniques (e.g., ReMix [6] and SmoothMix [13]) enhance model robustness, while more recent work introduces frequency- and structure-based modifications [11]. These augmentation strategies have proven highly effective in both supervised and self-supervised learning. In contrast to NLP and computer vision, data augmentation in PPM remains underexplored. The work in [12] introduced the first comprehensive model-agnostic augmentation framework for event logs. Their method applies a set of timestamp-preserving transformations, including random insertion, deletion, replacement and swap (derived from [22]), but also introduces swapping of events with the same timestamp and loop augmentation. Although these transformations increase variability, a key limitation is that they are partly applied at random, without considering the underlying process structure. This can result in unrealistic behavior that contradicts the actual control-flow semantics of the process. As such, while the method enriches the training data, it does not guarantee that the new traces are useful or even likely within the process domain. This gap motivates the development of statistically grounded transformations, which are transformations based on frequent paths or XOR-splits mined from the original training data. Such augmentations can maintain control-flow while still introducing meaningful variability to support generalization.
@@ 2.0.2 Self-Supervised Learning (SSL) Approaches
have emerged as powerful tools to learn useful representations from unlabeled data. Instead of relying on manual annotations, SSL defines pretext tasks that use the data itself to generate training input. This has proven particularly effective in domains where labeled data is scarce or expensive to obtain. A dominant class of SSL methods uses contrastive learning techniques. SimCLR [4] generates two augmented views of the same input and trains the model to bring these closer in the embedding space while pushing apart other (negative) samples. While effective, contrastive methods require careful design of negative pairs, large batch sizes and heavy data augmentation. This makes them difficult to apply to structured, small-scale domains such as PPM. To overcome these limitations, Siamese learning frameworks like BYOL [10] and SimSiam [5] have gained popularity (cf. Figure 2). These models use two identical networks to process different augmented versions of the same input, learning to align their latent representations. Unlike contrastive methods, they do not require negative pairs or large batch sizes. BYOL introduces a momentum-updated target network, while SimSiam claims to achieve similar performance with simpler architectures and a stop-gradient mechanism [5]. Both frameworks have shown strong results in learning high-quality embeddings across NLP and computer vision tasks. SSL enables representation learning from unlabeled data, which can benefit PPM tasks such as next activity prediction and final outcome prediction. Pre-training with SSL can also help when labeled traces are limited or imbalanced, which is often the case, especially for final outcome prediction. Moreover, SSL encourages the model to learn structural and semantic properties of the data, rather than overfitting to specific task labels. However, the quality of learned representations is clearly dependent on the diversity and the relevance of the augmentations used to create positive pairs. This makes statistically grounded data augmentation not just beneficial, but essential for effective representation learning in PPM.
@@ 3 Problem Definition
Let =$\mathcal{L}=\{\sigma_{1},\sigma_{2},\dots,\sigma_{N}\}$ denote an event log composed of $N$ traces, where each trace $\sigma_{i}=\langle e_{i,1},e_{i,2},\dots,e_{i,T_{i}}\rangle$ is a sequence of $T_{i}$ events associated with a process instance. Each event $e_{i,t}$ carries attributes such as the activity label, timestamp and possibly additional case or event-level features. Each event $e_{i,t}$ is here intended as a tuple $(A_{i,t},t_{i,t},\vec{f}_{i,t})$, in which the activity $A_{i,t}\in\mathcal{A}$ for $\mathcal{A}$ the set of possible activities of the event log $\mathcal{L}$, $t_{i,t}\in\mathcal{T}$ the set of admitted timestamps and $\vec{f}\in\mathcal{F}_{1}\times\ldots\times\mathcal{F}_{m}$, where $\mathcal{F}_{1},\ldots,\mathcal{F}_{m}$ are the sets containing the values of the different features. To lighten the notation, in this paper we are going to refer to $\langle A_{1},\ldots,A_{n}\rangle$ as the sequence of events $\langle e_{1},\ldots,e_{n}\rangle$ for which the corresponding activities are $A_{1},\ldots,A_{n}$ and both the timestamps and the feature vectors are not known. Now, it is possible to define the prediction problems as follows:
@@ 4.1 Novel Transformation Methods
To keep the semantics of process traces and effectiveness of data augmentation in PPM, we propose three statistically grounded transformation techniques, named StatisticalInsertion, StatisticalDeletion and StatisticalReplacement (Fig. 3). Unlike traditional random transformations, these methods leverage frequent activity patterns in the event log to generate realistic and process-compliant trace variants. The techniques are parameterized by thresholds $\alpha,\beta,\gamma,\delta\in[0,1]$ and a maximum intermediate length $\lambda_{\max}$, which control frequency constraints and structural complexity.
@@ 4.2 Data Augmentation Strategy
To generate diverse yet semantically coherent process trace variants, we apply two successive augmentations to each training prefix using a pool of transformation functions $\mathcal{Z}$. Our pipeline consists of two core stages: i) applicability filtering and a ii) pairwise augmentation check. Given a prefix of a trace $x$, we first determine which augmentors are applicable by evaluating each method’s structural constraints. Our main augmentors (StatisticalInsertion, StatisticalDeletion and StatisticalReplacement) are applied when the sequence matches known behavioral patterns mined from the training log (Section 4.1). If none are applicable, we fall back to simpler, structure-agnostic augmentors named RandomInsertion, RandomDeletion and RandomReplacement. For each input prefix $x$, we record all augmentors that are valid candidates for transformation. Next, we generate two augmented views $v=t(x),v^{\prime}=t^{\prime}(x)$, where $t,t^{\prime}\in\mathcal{Z}$ are transformation functions sampled from the augmentation pool $\mathcal{Z}$, such that $v\neq v^{\prime}$. We first sample an applicable augmentor uniformly and apply it to produce $v$. Then, we attempt up to 30 trials to generate a second distinct augmentation $v^{\prime}$, avoiding trivial duplicates. Each augmented pair is padded using left-padding to a fixed length determined by the longest transformed sequence in the batch. This ensures input consistency for self-supervised learning. This strategy ensures each trace contributes semantically meaningful variation to the training set, leveraging both structural knowledge (from frequent patterns) and generalization capability (via fallback augmentors).
@@ 4.3 Siamese Pre-Training
We symmetrize the loss $\mathcal{L}_{\theta,\xi}$ in Equation 3 by separately feeding $v$ to the online network and $v^{\prime}$ to the target network to compute both $\mathcal{L}_{\theta,\xi}$ and ~$\tilde{\mathcal{L}}_{\theta,\xi}$. The final training objective is the symmetric loss:
@@
where $\tau\in[0,1)$ is the momentum coefficient. This setup ensures that the models learns that the model learns to align the online representation with a slowly evolving target representation, allowing robust and stable learning even in the absence of negative samples.
@@ 5.0.4 Implementation Details
The implementation including the model parameters is available on GitHub  [^3] https://github.com/SvStraten/SiamSA-PPM. Experiments have been performed in Python 3.12.2 using TensorFlow 2.15. We used a temporal train-validation-test split of 65-15-20. To ensure fair comparison, each model is trained and evaluated over five independent repetitions, and we did not run any experiments in parallel. We performed sensitivity analysis on the BPIC13-c and BPIC13-i datasets. After testing values in the range $10^{-2}$ to $10^{-5}$, we selected $10^{-4}$ as the optimal value for $\alpha$, $\beta$, $\gamma$, and $\delta$. For $\lambda_{\text{max}}$, we have tested integers between 2 and 5, after which 4 has been chosen as the optimal value.s
@@ 6.0.2 Effect of Augmentation on Performance
We assess the impact of event log augmentation on next activity prediction across five datasets: Sepsis, BPIC15-1, BPIC13-i, BPIC13-c, and Helpdesk. Each baseline model is evaluated on the original (base) logs, logs augmented using the method in [12] and logs augmented with our proposed strategy (Section 4.2). Table 1 reports results using the best-performing augmentation factor (1.2, 1.5 or 2.0). On BPIC15-1, our method increases performance for all baselines, while the augmentation strategy in [12] does not consistently yield improvements. For the other datasets, both methods typically lead to gains over the base logs, with most models showing higher accuracy when trained on augmented data. Notably, on Sepsis and BPIC13-i almost all baselines benefit from augmentation. These results confirm that augmentation, whether via our approach or prior work, can enhance predictive performance, though the effectiveness varies by dataset and model.
Table 1: Accuracy of the baselines for next activity prediction using base event log, the augmented log following [12] and the augmented log following our method (Section 4.2).
@@ 6.0.4 Final Outcome Prediction
Table 3 summarizes the accuracy results for binary classification targets on the BPIC12 and Sepsis datasets on Final Outcome Prediction problem (cf. Section 3). It is important to note that not all datasets are equally relevant for this task, as some datasets do not include relevant outcome targets. BPIC13-c BPIC13-i and Helpdesk do not contain varying outcomes for instance. For BPIC12, the outcomes correspond to the status of financial application requests: Approved, Declined and Cancelled. Our model consistently ranks near the top on all three targets, reaching 80.04% accuracy on the Declined class, just 0.82% below the top performer and closely matches the best scores for Approved and Cancelled. In the Sepsis dataset, the prediction targets indicate the reason for patient release after treatment. These include Release-A, Release-B, Release-C and Release-D, which reflect different clinical justifications for discharge. SiamSA-PPM achieves the highest accuracy for Release-B (94.21%) and Release-D (96.02%), and remains competitive on the remaining targets, demonstrating robust performance in distinct clinical outcomes.

### v2

@@ 1 Introduction
Due to the increasing digitization of business operations, Predictive Process Monitoring (PPM) has become a critical area of research within process mining. Two of its core tasks, next activity prediction and final outcome prediction, have received substantial attention due to their practical relevance in optimizing workflows, reducing delays, and improving resource utilization [20]. While earlier works utilized symbolic models such as Petri nets and transition systems, recent advancements in deep learning, especially sequence models such as Long-Short Term Memory (LSTM), Gated Recurrent Unit and attention-based Transformers, have significantly improved the predictive performance of such models. However, a persistent challenge in deep learning and inherently in its application to PPM, is the limited size of available labeled data [3]. Despite their high complexity, business processes are typically logged with limited trace variation and often result in datasets containing redundant or highly similar samples. These limitations affect the generalizability of predictive models, especially in scenarios with infrequent behaviors or class imbalance. In machine learning, a common remedy for such constraints is data augmentation: the process of artificially generating additional training data to enhance diversity and robustness. Data augmentation has been extensively explored in domains such as computer vision and natural language processing (NLP) [21]. In these fields, augmentations are informed by domain knowledge to preserve the semantic integrity of the data. For example, visual enhancements such as flipping or cropping retain object identity, while textual methods such as synonym replacement, word deletion or insertion aim to maintain grammatical and contextual coherence [22]. The core principle in these domains is that augmentations should be semantically valid and task-relevant. In contrast, data augmentation for process mining remains considerably underexplored. Prior work explored model-agnostic augmentation using primarily random insertions, deletions or replacements [12], but such random transformations risk violating process semantics. For example, delivering an order before it has been shipped (cf. Fig. 1). In contrast, our approach introduces statistically grounded augmentation techniques that leverage frequent control-flow patterns to generate realistic, semantically valid trace variants, thereby aligning augmentation with actual process dynamics (cf. Section 2). We use the process in Fig. 1 $(Order\rightarrow Pack\rightarrow Ship\rightarrow Deliver)$ as a running example throughout the paper to illustrate how random augmentations can violate semantics, while our statistically grounded methods preserve realistic control-flow.
@@
In addition to augmentation, we also explore the underutilized potential of self-supervised learning (SSL) in PPM. While most existing approaches rely on supervised learning, this requires labeled data (e.g. next activity or final outcome annotations) for each trace. Practically, such amount of labels may not be sufficient to train a model. In fact, some event logs may be imbalanced, resulting in, for example, an insufficient number of examples for a given outcome, e.g. a given machine breaks down. This issue also applies to the next activity prediction task, where the occurrence of certain activities may be rare, leading to a very limited number of events with the corresponding label. SSL addresses this by learning from the data itself without requiring manual labels. It constructs surrogate tasks, such as aligning different augmented views of the same input to learn robust task-agnostic representations. Although SSL has seen great success in the computer vision domain [4, 10, 5], it is underexplored in PPM. In this work, we adopt a Siamese learning strategy based on Bootstrap Your Own Latent (BYOL) [10], where augmented trace prefixes are encoded to similar latent representations. This setup avoids the need for negative samples or large batch sizes [10], making it well-suited for structured, small-scale process logs. Combined with our statistically grounded augmentations, this enables high-quality representation for downstream tasks, even with limited labeled data.
@@ 2.0.1 Data Augmentation Techniques
Data augmentation is a proven strategy to increase training data and improve model generalization. In NLP, a well-known method is EDA [22], which introduces random augmentations such as random insertion, deletion and replacement of words. Other techniques such as syntax-aware transformations [8] and generative approaches such as InsNet [15] inject linguistic diversity into text classification and translation tasks. In computer vision, simple pixel-level augmentations and mixing techniques (e.g. ReMix [6] and SmoothMix [13]) enhance model robustness, while more recent work introduces frequency- and structure-based modifications [11]. These augmentation strategies have proven highly effective in both supervised and self-supervised learning. In contrast to NLP and computer vision, data augmentation in PPM remains underexplored. The work in [12] introduced the first comprehensive model-agnostic augmentation framework for event logs. Their method applies a set of timestamp-preserving transformations, including random insertion, deletion, replacement and swap (derived from [22]), but also introduces swapping of events with the same timestamp and loop augmentation. Although these transformations increase variability, a key limitation is that they are partly applied at random, without considering the underlying process structure. This can result in unrealistic behavior that contradicts the actual control-flow semantics of the process. As such, while the method enriches the training data, it does not guarantee that the new traces are useful or even likely within the process domain. This gap motivates the development of statistically grounded transformations, which are transformations based on frequent paths or XOR-splits mined from the original training data. Such augmentations can maintain control-flow while still introducing meaningful variability to support generalization.
@@ 2.0.2 Self-Supervised Learning (SSL) Approaches
have emerged as powerful tools to learn useful representations from unlabeled data. Instead of relying on manual annotations, SSL defines pretext tasks that use the data itself to generate training input. This has proven particularly effective in domains where labeled data is scarce or expensive to obtain. A dominant class of SSL methods uses contrastive learning techniques. SimCLR [4] generates two augmented views of the same input and trains the model to bring these closer in the embedding space while pushing apart other (negative) samples. While effective, contrastive methods require careful design of negative pairs, large batch sizes and heavy data augmentation. This makes them difficult to apply to structured, small-scale domains such as PPM. To overcome these limitations, Siamese learning frameworks like BYOL [10] and SimSiam [5] have gained popularity (cf. Fig. 2). These models use two identical networks to process different augmented versions of the same input, learning to align their latent representations. Unlike contrastive methods, they do not require negative pairs or large batch sizes. BYOL introduces a momentum-updated target network, while SimSiam claims to achieve similar performance with simpler architectures and a stop-gradient mechanism [5]. Both frameworks have shown strong results in learning high-quality embeddings across NLP and computer vision tasks.
SSL enables representation learning from unlabeled data, which can benefit PPM tasks such as next activity prediction and final outcome prediction. Pre-training with SSL can also help when labeled traces are limited or imbalanced, which is often the case, especially for final outcome prediction. Moreover, SSL encourages the model to learn structural and semantic properties of the data, rather than overfitting to specific task labels. However, the quality of learned representations is clearly dependent on the diversity and the relevance of the augmentations used to create positive pairs. This makes statistically grounded data augmentation not just beneficial, but essential for effective representation learning in PPM.
@@ 3 Problem Definition
Let $\mathcal{L}=\{\sigma_{1},\sigma_{2},\dots,\sigma_{N}\}$ denote an event log composed of $N$ traces, where each trace $\sigma_{i}=\langle e_{i,1},e_{i,2},\dots,e_{i,T_{i}}\rangle$ is a sequence of $T_{i}$ events associated with a process instance. Each event $e_{i,t}$ carries attributes such as the activity label, timestamp and possibly additional case or event-level features. Each event $e_{i,t}$ is here intended as a tuple $(A_{i,t},t_{i,t},\vec{f}_{i,t})$, in which the activity $A_{i,t}\in\mathcal{A}$ for $\mathcal{A}$ the set of possible activities of the event log $\mathcal{L}$, $t_{i,t}\in\mathcal{T}$ the set of admitted timestamps and $\vec{f}\in\mathcal{F}_{1}\times\ldots\times\mathcal{F}_{m}$, where $\mathcal{F}_{1},\ldots,\mathcal{F}_{m}$ are the sets containing the values of the different features. To lighten the notation, in this paper we are going to refer to $\langle A_{1},\ldots,A_{n}\rangle$ as the sequence of events $\langle e_{1},\ldots,e_{n}\rangle$ for which the corresponding activities are $A_{1},\ldots,A_{n}$ and both the timestamps and the feature vectors are not known. In line with common practice in PPM, we treat each recorded event as the completion of an activity. Now, it is possible to define the prediction problems as follows:
@@ 4.1 Novel Transformation Methods
To keep the semantics of process traces and effectiveness of data augmentation in PPM, we propose three statistically grounded transformation techniques, named StatisticalInsertion, StatisticalDeletion and StatisticalReplacement (cf. Fig. 3). Unlike traditional random transformations, these methods leverage frequent activity patterns in the event log to generate realistic and process-compliant trace variants. The techniques are parameterized by thresholds $\alpha,\beta,\gamma,\delta\in[0,1]$ and a maximum intermediate length $\lambda_{\max}$, which control frequency constraints and structural complexity.
@@
For instance, in our running example, StatisticalInsertion might add a frequent step $CheckDocs$ between $Order$ and $Pack$, while StatisticalReplacement could substitute one observed path $Pack\rightarrow Ship\rightarrow Deliver$ with a special or quick delivery path $Pack\rightarrow ExpressShip\rightarrow Deliver$.
@@ 4.2 Data Augmentation Strategy
To generate diverse yet semantically coherent process trace variants, we apply two successive augmentations to each training prefix using a pool of transformation functions $\mathcal{Z}$. Our pipeline consists of two core stages: i) applicability filtering and a ii) pairwise augmentation check. Given a prefix of a trace $x$, we first determine which augmentors are applicable by evaluating each method’s structural constraints. Our main augmentors (StatisticalInsertion, StatisticalDeletion and StatisticalReplacement) are applied when the sequence matches known behavioral patterns mined from the training log (cf. Section 4.1). If none are applicable, we fall back to simpler, structure-agnostic augmentors named RandomInsertion, RandomDeletion and RandomReplacement. For each input prefix $x$, we record all augmentors that are valid candidates for transformation. Next, we generate two augmented views $v=t(x),v^{\prime}=t^{\prime}(x)$, where $t,t^{\prime}\in\mathcal{Z}$ are transformation functions sampled from the augmentation pool $\mathcal{Z}$, such that $v\neq v^{\prime}$. We first sample an applicable augmentor uniformly and apply it to produce $v$. Then, we attempt up to 30 trials to generate a second distinct augmentation $v^{\prime}$, avoiding trivial duplicates. Each augmented pair is padded using left-padding to a fixed length determined by the longest transformed sequence in the batch. This ensures input consistency for self-supervised learning. This strategy ensures each trace contributes semantically meaningful variation to the training set, leveraging both structural knowledge (from frequent patterns) and generalization capability (via fallback augmentors).
For instance, in our running example, the prefix $Order\rightarrow Pack$ allows only insertion as valid transformation. Among the eligible candidates (derived from StatisticalInsertion), both $CheckDocs$ and $ValidatePayment$ can be inserted, yielding $Order\rightarrow CheckDocs\rightarrow Pack$ and $Order\rightarrow ValidatePayment\rightarrow Pack$ as the two augmented prefixes. This illustrates how the strategy selects multiple valid options to generate distinct yet semantically consistent training views.
@@ 4.3 Siamese Pre-Training
We symmetrize the loss $\mathcal{L}_{\theta,\xi}$ in Eq. 3 by separately feeding $v$ to the online network and $v^{\prime}$ to the target network to compute both $\mathcal{L}_{\theta,\xi}$ and $\tilde{\mathcal{L}}_{\theta,\xi}$. The final training objective is the symmetric loss:
@@
where $\tau\in[0,1)$ is the momentum coefficient. This setup ensures that the models learns to align the online representation with a slowly evolving target representation, allowing robust and stable learning even in the absence of negative samples.
@@ 5.0.4 Implementation Details
The implementation including the model parameters is available on GitHub  [^3] https://github.com/SvStraten/SiamSA-PPM. Experiments have been performed in Python 3.12.2 using TensorFlow 2.15. We used a temporal train-validation-test split of 65-15-20. To ensure fair comparison, each model is trained and evaluated over five independent repetitions, and we did not run any experiments in parallel. We performed sensitivity analysis on the BPIC13-c and BPIC13-i datasets. After testing values in the range $10^{-2}$ to $10^{-5}$, we selected $10^{-4}$ as the optimal value for $\alpha$, $\beta$, $\gamma$, and $\delta$. For $\lambda_{\text{max}}$, we have tested integers between 2 and 5, after which 4 has been chosen as the optimal value.
@@ 6.0.2 Effect of Augmentation on Performance
We assess the impact of event log augmentation on next activity prediction across five datasets: Sepsis, BPIC15-1, BPIC13-i, BPIC13-c, and Helpdesk. Each baseline model is evaluated on the original (base) logs, logs augmented using the method in [12] and logs augmented with our proposed strategy (cf. Section 4.2). Table 1 reports results using the best-performing augmentation factor (1.2, 1.5 or 2.0). On BPIC15-1, our method increases performance for all baselines, while the augmentation strategy in [12] does not consistently yield improvements. For the other datasets, both methods typically lead to gains over the base logs, with most models showing higher accuracy when trained on augmented data. Notably, on Sepsis and BPIC13-i almost all baselines benefit from augmentation. These results confirm that augmentation, whether via our approach or prior work, can enhance predictive performance, though the effectiveness varies by dataset and model.
Table 1: Accuracy of the baselines for next activity prediction using base event log, the augmented log following [12] and the augmented log following our method (cf. Section 4.2).
@@ 6.0.4 Final Outcome Prediction
Table 3 summarizes the accuracy results for binary classification targets on the BPIC12 and Sepsis datasets on Final Outcome Prediction problem (cf. Section 3). It is important to note that not all datasets are equally relevant for this task, as some datasets do not include relevant outcome targets. BPIC13-c, BPIC13-i and Helpdesk do not contain varying outcomes for instance. For BPIC12, the outcomes correspond to the status of financial application requests: Approved, Declined and Cancelled. Our model consistently ranks near the top on all three targets, reaching 80.04% accuracy on the Declined class, just 0.82% below the top performer and closely matches the best scores for Approved and Cancelled. In the Sepsis dataset, the prediction targets indicate the reason for patient release after treatment. These include Release-A, Release-B, Release-C and Release-D, which reflect different clinical justifications for discharge. SiamSA-PPM achieves the highest accuracy for Release-B (94.21%) and Release-D (96.02%), and remains competitive on the remaining targets, demonstrating robust performance in distinct clinical outcomes.